# Intention Analysis Makes LLMs A Good Jailbreak Defender

### v3

@@ Abstract
Aligning large language models (LLMs) with human values, particularly in the face of complex and stealthy jailbreak attacks, presents a formidable challenge. In this study, we present a simple yet highly effective defense strategy, i.e., Intention Analysis ($\mathbb{IA}$). The principle behind this is to trigger LLMs’ inherent self-correct and improve ability through a two-stage process: 1) essential intention analysis, and 2) policy-aligned response. Notably, $\mathbb{IA}$ is an inference-only method, thus could enhance the safety of LLMs without compromising their helpfulness [^1] Improving the safety of LLM with training-required methods Ouyang et al. (2022); Touvron et al. (2023) always necessitates great effort to strike a delicate balance between safety and helpfulness.. Extensive experiments on varying jailbreak benchmarks across ChatGLM, LLaMA2, Vicuna, MPT, DeepSeek, and GPT-3.5 show that $\mathbb{IA}$ could consistently and significantly reduce the harmfulness in responses (averagely -53.1% attack success rate) and maintain the general helpfulness. Encouragingly, with the help of our $\mathbb{IA}$, Vicuna-7B even outperforms GPT-3.5 in terms of attack success rate. Further analyses present some insights into how our method works. To facilitate reproducibility, we release our code and scripts at: https://github.com/alphadl/SafeLLM_with_IntentionAnalysis.
@@
Yuqi Zhang1, Liang Ding2, Lefei Zhang1 Dacheng Tao2 1School of Computer Science, Wuhan University 2School of Computer Science, The University of Sydney {zhangyuqi, zhanglefei}@whu.edu.cn, {liangding.liam, dacheng.tao}@gmail.com
@@ 1 Introduction
Recently, Large Language Models (LLMs) (Touvron et al., 2023; OpenAI, 2023; Google, 2023), such as ChatGPT, not only show remarkable capabilities in various tasks Qin et al. (2023); Zhong et al. (2023); Peng et al. (2023), but also lead to the risk of potential misuse (e.g., producing harmful responses or illegal suggestions) Weidinger et al. (2021). Efforts like Reinforcement Learning from Human Feedback (RLHF, Ouyang et al., 2022) have been made to alleviate these risks and enhance LLMs’ alignment with human values, making LLMs able to refuse harmful questions like how to rob a bank? However, LLMs remain vulnerable to some adversarial inputs, particularly in the context of so-called “jailbreak” attacks. These jailbreak attacks are specially designed to circumvent the safety policy and manipulate LLMs for their restricted outputs (Yuan et al., 2024; Zou et al., 2023), which poses formidable risks in real applications.
@@
Despite numerous methods focusing on defending LLMs against these jailbreak attacks (Li et al., 2024; Jain et al., 2023; Chen et al., 2024), their efficacy markedly decreases when handling those with stealthy and complex intention (see experimental results in Section 4.2). This type of jailbreak attacks are typically hand-crafted prompts by using more instructions and employing specific techniques, such as creating a virtual scenario where LLMs are assigned the role of “DAN” (Do Anything Now, Liu et al., 2023b). Such complex and stealthy intents pose a great challenge for LLMs to maintain safety when confronting these jailbreak attacks (Shen et al., 2023; Bao et al., 2023).
In this paper, we leverage the intrinsic intent recognition capabilities of LLMs, proposing an Intention Analysis ($\mathbb{IA}$) strategy, which cleverly addresses the aforementioned challenge. $\mathbb{IA}$ enables LLMs to recognize the underlying intention of the user query to better understand it and perceive the unsafe content within before responding, therefore significantly enhancing their safety against varying jailbreak attacks, see Figure 1 for a demonstration. Practically, $\mathbb{IA}$ follows a two-stage process, where $\mathbb{IA}$ first directs LLMs to analyze the essential intention of the user query, and then extract the safety policy-aligned final response from LLMs, as shown in Figure 2. Notably, our $\mathbb{IA}$ is an inference-only method that can significantly enhance LLM safety without the need for additional safety training (Ouyang et al., 2022; Touvron et al., 2023). In this way, $\mathbb{IA}$ skillfully circumvents the safety-helpfulness trade-off and enables comparable safety improvement as well as better helpfulness maintenance.
@@ 2 Related Work
Despite significant efforts to align LLMs with human preference (Ouyang et al., 2022; Bai et al., 2022; Lee et al., 2023; Korbak et al., 2023; Miao et al., 2024), adversarial attackers can still elicit harmful responses from LLMs by “jailbreak” attacks (Shen et al., 2023; Liu et al., 2023b). Current jailbreak attack methods are primarily classified into two categories: in-the-wild jailbreak prompts and optimization-based automatic attacks (Chao et al., 2023). In-the-wild jailbreak prompts are typically hand-crafted through human ingenuity and is semantically understandable in general Shen et al. (2023). For optimization-based automatic attacks, a representative work is to automatically fetch a transferable attack suffix through the Greedy Coordinate Gradient (GCG) algorithm which maximizes the probability of the language model generating an affirmative and unsafe response Zou et al. (2023).
@@
Recently, numerous methods have been developed to reduce LLMs’ harmful generations in inference stage. A branch of them mainly concentrates on controlling the content that LLMs can see by pre-processing the user inputs, such as perplexity filtering (Alon and Kamfonas, 2023; Jain et al., 2023), paraphrasing (Jain et al., 2023) and re-tokenization (Cao et al., 2023; Jain et al., 2023). Another branch focuses on exploiting LLMs’ intrinsic capabilities of self-correction and improvement against jailbreak attacks, such as letting LLMs self-evaluate their outputs (Helbling et al., 2023) or reminding of safety in system mode with conventional decoding Xie et al. (2023) or contrastive decoding Zhong et al. (2024).
While these existing methods show effectiveness in preventing LLMs from generating unsafe responses, their efficacy significantly diminishes when confronted with sophisticated jailbreak attacks due to the ignorance of the underlying stealthy and malicious intention. In contrast, our method handles such jailbreak attacks by leveraging the intrinsic intent recognition capabilities of LLMs, which leads to a significant enhancement of LLM safety (see Table 1 for details).
@@ 3.1 Preliminary
$q(y_{1:L}|x_{1:n}^{s},x_{1:m}^{u})=\prod\limits_{i=1}^{L}q(y_{i}|y_{1:i-1},x_{1:n}^{s},x_{1:m}^{u}).$
For simplicity, we write $R\sim q(R|P_{sys},P_{usr)}$ to denote sampling a response $R$ from $q(\cdot)$ given the prompt $P_{sys}$ and $P_{usr}$. In this way, the response $R$ can be obtained as: $R=\textrm{LLM}\left(P_{sys},P_{usr}\right).$
@@ 3.2 $\mathbb{IA}$: Intention Analysis
This stage focuses on guiding the LLMs to discern the core intention behind the user query, with a specific orientation towards safety, ethics, and legality. The critical question arises: How can we ensure that LLMs accurately identify the query’s intention? Actually, recent studies Bender and Koller (2020); Zhu et al. (2024); Gómez-Pérez et al. (2023) have shown that LLMs are notably proficient at language understanding tasks, and intention analysis is a straightforward task, indicating the competence of LLMs in performing this stage. The only concern is generative models’ potential hallucination when performing the discriminative tasks Ji et al. (2023); Yan et al. (2021); Ye et al. (2023), therefore, we carefully define the format for the models’ response, that is, beginning with “The essential intention of the query is”, which has been validated in our analysis.
@@
$R_{st1}=\textrm{LLM}\left(P_{sys},I_{rec}\oplus P_{usr}\right),$
@@
Having successfully recognized the essential intention, the second stage aims to elicit the final response from the LLMs. We first direct the LLMs to bear the identified intention in mind and then provide a final response to the user query. Meanwhile, we explicitly instruct the LLMs to strictly adhere to safety policy and ethical standards and ensure the exclusion of any unsafe content in their responses. To this end, the second stage further strengthens the role of the intention analysis and reinforces the inherent alignment of LLMs with the safety policy.
@@
$R_{st2}=\textrm{LLM}\left(P_{sys},I_{rec}\oplus P_{usr},R_{st1},I_{ct}\right).$
@@
Models Defense Methods Jailbreak Methods Average
@@
ChatGLM-6B Vanilla 29.0 45.8 100 88.0 99.5 72.5
ICD 19.1 2.81 17.1 17.0 2.00 11.6
Self-Reminder 22.5 3.13 17.9 0.00 66.0 21.9
SmoothLLM 7.19 20.6 84.5 1.00 84.0 39.5
$\mathbb{IA}$ (Ours) 5.48 6.12 0.00 1.00 2.00 2.92
@@
ICD 0.98 0.00 0.00 0.00 0.00 0.20
Self-Reminder 0.77 0.00 4.38 0.00 0.00 1.03
SmoothLLM 0.31 2.81 86.5 0.00 71.5 32.2
$\mathbb{IA}$ (Ours) 0.13 0.00 0.00 0.00 0.00 0.03
@@
ICD 40.4 32.8 0.00 1.00 88.0 32.4
Self-Reminder 41.3 33.8 55.4 11.0 98.5 48.0
SmoothLLM 13.5 54.4 96.4 8.00 98.5 54.2
$\mathbb{IA}$ (Ours) 3.42 0.31 0.00 0.00 10.5 2.85
@@
ICD 53.9 32.8 86.9 0.00 91.5 53.0
Self-Reminder 52.5 36.9 75.7 1.00 83.0 49.8
SmoothLLM 17.3 37.0 94.0 5.00 98.0 50.3
$\mathbb{IA}$ (Ours) 0.94 1.12 0.00 0.00 3.50 1.11
@@
ICD 49.4 29.9 100 3.00 45.6
Self-Reminder 46.9 39.4 100 19.0 51.3
SmoothLLM 60.6 64.4 22.0 22.0 42.3
$\mathbb{IA}$ (Ours) 5.38 19.2 4.78 4.00 8.34
@@
ICD 45.6 14.4 47.8 9.00 29.2
Self-Reminder 9.58 7.81 3.19 1.00 5.40
SmoothLLM 26.9 11.9 51.0 0.00 22.4
$\mathbb{IA}$ (Ours) 3.78 1.56 7.57 2.00 3.73
@@
ICD 0.94 0.31 0.00 0.00 0.31
Self-Reminder 2.81 0.31 0.00 0.00 0.78
SmoothLLM 0.64 0.00 0.00 0.00 0.16
$\mathbb{IA}$ (Ours) 0.64 0.00 0.00 0.00 0.16
Table 1: Comparison of our $\mathbb{IA}$ and four baselines under five jailbreak methods in terms of ASR (%). The best average results are highlighted in bold. Among them, DAN, SAP200, and DeepInception are complex and stealthy in-the-wild jailbreaks, while GCG and AutoDAN are optimization-based automatic jailbreaks. “—” means lacking official AutoDAN implementation for distributed larger models (MPT-30B and DeekSeek-67B) or white-box LLM weights required (GPT-3.5).

### v4

@@ Abstract
Aligning large language models (LLMs) with human values, particularly when facing complex and stealthy jailbreak attacks, presents a formidable challenge. Unfortunately, existing methods often overlook this intrinsic nature of jailbreaks, which limits their effectiveness in such complex scenarios. In this study, we present a simple yet highly effective defense strategy, i.e., Intention Analysis ($\mathbb{IA}$). $\mathbb{IA}$ works by triggering LLMs’ inherent self-correct and improve ability through a two-stage process: 1) analyzing the essential intention of the user input, and 2) providing final policy-aligned responses based on the first round conversation. Notably, $\mathbb{IA}$ is an inference-only method, thus could enhance LLM safety without compromising their helpfulness [^1] Improving the safety of LLM with training-required methods Ouyang et al. (2022); Touvron et al. (2023) always necessitates great effort to strike a delicate balance between safety and helpfulness.. Extensive experiments on varying jailbreak benchmarks across a wide range of LLMs show that $\mathbb{IA}$ could consistently and significantly reduce the harmfulness in responses (averagely -48.2% attack success rate). Encouragingly, with our $\mathbb{IA}$, Vicuna-7B even outperforms GPT-3.5 regarding attack success rate. We empirically demonstrate that, to some extent, $\mathbb{IA}$ is robust to errors in generated intentions. Further analyses reveal the underlying principle of $\mathbb{IA}$: suppressing LLM’s tendency to follow jailbreak prompts, thereby enhancing safety.
@@
Yuqi Zhang1, Liang Ding2, Lefei Zhang1thanks:  Corresponding Author., Dacheng Tao3 1School of Computer Science, Wuhan University   2The University of Sydney 3College of Computing and Data Science at Nanyang Technological University, Singapore 639798 [Uncaptioned image] {zhangyuqi,zhanglefei}@whu.edu.cn,  {liangding.liam,dacheng.tao}@gmail.com [Uncaptioned image] https://github.com/alphadl/SafeLLM_with_IntentionAnalysis
@@ 1 Introduction
Recently, Large Language Models (LLMs) (Touvron et al., 2023; OpenAI, 2023; Google, 2023), such as ChatGPT, not only show remarkable capabilities in various tasks Qin et al. (2023); Zhong et al. (2023); Peng et al. (2023); Ren et al. (2024), but also lead to the risk of potential misuse (e.g., producing harmful responses or illegal suggestions) Weidinger et al. (2021). Efforts like Reinforcement Learning from Human Feedback (RLHF, Ouyang et al., 2022) have been made to alleviate these risks and enhance LLMs’ alignment with human values, making LLMs able to refuse direct harmful questions like how to rob a bank? However, LLMs remain vulnerable to some adversarial inputs, particularly in the context of so-called “jailbreak” attacks. These jailbreaks are specially designed to circumvent safety policy and manipulate LLMs for their restricted outputs (Yuan et al., 2024; Zou et al., 2023), which poses formidable risks in real applications.
@@
To defend LLMs against jailbreak attacks, existing popular methods either focus on emphasizing safety during inference (Xie et al., 2023; Wei et al., 2023b), or modifying the user inputs (Robey et al., 2023) or evaluating inputs/outputs’ safety (Li et al., 2024), often neglecting the intrinsic characteristics of jailbreak attacks. This oversight limits their effectiveness in more complex jailbreak scenarios (see experimental results in Section 4.2). Through analysis, we find that these jailbreaks typically work by concealing harmful questions within seemingly inoffensive and complex scenarios, such as role-playing or virtual scene construction (Liu et al., 2023b). Such disguise leads LLMs to focus on the jailbreak prompt excessively, impairing their awareness of the harmful question itself (See Figure 5 for evidence). We assume such insufficient awareness of the harmful content concealed in complex jailbreak queries is the fundamental reason for LLM’s vulnerability to these attacks. Drawing insights from classic dialogue system design (Allen and Perrault, 1980), an effective solution is to tailor an intent recognition mechanism specifically for jailbreak scenarios to enhance LLM’s understanding of user queries regarding safety and improve its ability to recognize concealed harmful questions.
In this paper, we leverage the intrinsic intent recognition capabilities of LLMs, proposing an Intention Analysis ($\mathbb{IA}$) strategy. Specifically, $\mathbb{IA}$ enables LLMs to analyze the essential intention of the user query to better understand it and recognize the underlying unsafe content within before finally responding, as shown in Figure 2. Such intention analysis mechanism can significantly improve LLM safety against varying jailbreak attacks, see Figure 1 for a demonstration. We dive deeper from the perspective of attention scores and find that the underlying principle of $\mathbb{IA}$ is to suppress LLM’s tendency to follow jailbreak prompts. Notably, our $\mathbb{IA}$ is an inference-only method that can significantly enhance LLM safety without the need for additional safety training (Ouyang et al., 2022; Touvron et al., 2023). In this way, $\mathbb{IA}$ skillfully circumvents the safety-helpfulness trade-off and enables comparable safety improvement as well as better helpfulness maintenance.
@@ 2 Related Work
Despite significant efforts to align LLMs with human preference (Ouyang et al., 2022; Bai et al., 2022; Lee et al., 2023; Korbak et al., 2023; Miao et al., 2024), adversarial attackers can still elicit harmful responses from LLMs by “jailbreak” attacks (Shen et al., 2023; Liu et al., 2023b). Current jailbreak attack methods are primarily classified into two categories: in-the-wild jailbreak prompts and optimization-based automatic attacks (Chao et al., 2023; Yu et al., 2023). In-the-wild jailbreak prompts are typically hand-crafted through human ingenuity and is semantically understandable in general Shen et al. (2023). For optimization-based automatic attacks, a representative work is to automatically fetch a transferable attack suffix through the Greedy Coordinate Gradient (GCG) algorithm which maximizes the probability of the language model generating an affirmative and unsafe response Zou et al. (2023). In this work, various attacks mentioned above are considered in experiments to comprehensively test the defensive performance of our method.
@@
Recently, numerous methods have been developed to reduce LLMs’ harmful generations in inference stage. A branch of them mainly concentrates on controlling the content that LLMs can see by pre-processing user inputs, such as perplexity filtering (Alon and Kamfonas, 2023; Jain et al., 2023), paraphrasing (Jain et al., 2023) and re-tokenization (Cao et al., 2023; Jain et al., 2023). Another branch focuses on exploiting LLMs’ intrinsic capabilities of self-correction and improvement against jailbreak attacks, such as letting LLMs self-evaluate their outputs (Helbling et al., 2023; Li et al., 2024; Wang et al., 2024) or reminding of safety in system mode with conventional decoding Xie et al. (2023) or contrastive decoding Zhong et al. (2024).
While existing methods effectively prevent unsafe responses, their efficacy drops significantly against sophisticated jailbreak attacks that conceal harmful questions within complex and seemingly inoffensive scenarios. In contrast, our method enhances LLM safety by leveraging the intrinsic intent recognition capabilities of LLMs to detect these concealed threats (see Table 1 for details).
@@ 3.1 Preliminary
$\vspace{-0.2cm}q(y_{1:L}|x_{1:n}^{s},x_{1:m}^{u})=\prod\limits_{i=1}^{L}q(y_{i}|y_{1:i-1},x_{1:n}^{s},x_{1:m}^{u}).\vspace{-0.2cm}$
For simplicity, we use $R\sim q(R|P_{sys},P_{usr)}$ to denote sampling a response $R$ from $q(\cdot)$ given the prompt $P_{sys}$ and $P_{usr}$. In this way, the response $R$ can be obtained as: $R=\textrm{LLM}\left(P_{sys},P_{usr}\right).$
@@ 3.2 $\mathbb{IA}$: Intention Analysis
This stage focuses on guiding the LLMs to discern the core intention behind the user query, with a specific orientation towards safety, ethics, and legality. The critical question arises: How can we ensure that LLMs accurately identify the query’s intention? Actually, recent studies Bender and Koller (2020); Zhu et al. (2024); Gómez-Pérez et al. (2023) have shown that LLMs are notably proficient at language understanding tasks, and intention analysis is a straightforward task, indicating the competence of LLMs in performing this stage. The only concern is generative models’ potential hallucination when performing the discriminative tasks Ji et al. (2023); Yan et al. (2021); Ye et al. (2023); Lu et al. (2024), therefore, we carefully define the format for the models’ response, that is, beginning with “The essential intention of the query is”, which has been validated in our analysis.
@@
$R_{st1}=\textrm{LLM}\left(P_{sys},I_{rec}\oplus P_{usr}\right),\vspace{-0.3cm}$
@@
Having successfully recognized the essential intention, the second stage aims to elicit the final response from the LLMs. We first direct the LLMs to bear the identified intention in mind and then provide a final response to the user query. Meanwhile, we explicitly instruct the LLMs to strictly adhere to safety policy and ethical standards [^4] The details of safety policy and ethical standards are not explicitly provided because we have found they significantly increase inference costs with minimal benefit. We believe that LLMs, through training, develop an inherent understanding of safety, allowing implicit prompts to effectively activate this internal knowledge. and ensure the exclusion of any unsafe content in their responses. To this end, the second stage further strengthens the role of the intention analysis and reinforces the inherent alignment of LLMs with the safety policy.
@@
$R_{st2}=\textrm{LLM}\left(P_{sys},I_{rec}\oplus P_{usr},R_{st1},I_{ct}\right).\vspace{-0.2cm}$
@@
Models Defense Methods Attack Methods Average Time Cost
@@
ChatGLM-6B Vanilla 29.0 45.8 100 88.0 99.5 72.5 14.3
Input Check 16.3 9.52 46.2 9.00 89.0 34.0 12.6
ICD 19.1 2.81 17.1 17.0 2.00 11.6 15.2
Self-Reminder 22.5 3.13 17.9 0.00 66.0 21.9 17.1
SmoothLLM 7.19 20.6 84.5 1.00 84.0 39.5 113.4
$\mathbb{IA}$ (Ours) 5.48 6.12 0.00 1.00 2.00 2.92 19.2
@@
Input Check 7.50 0.00 0.00 0.00 43.0 10.1 10.7
ICD 0.98 0.00 0.00 0.00 0.00 0.20 15.5
Self-Reminder 0.77 0.00 4.38 0.00 0.00 1.03 14.8
SmoothLLM 0.31 2.81 86.5 0.00 71.5 32.2 118.5
$\mathbb{IA}$ (Ours) 0.13 0.00 0.00 0.00 0.00 0.03 19.5
@@
Input Check 3.43 0.00 7.57 0.00 7.00 3.60 4.98
ICD 0.63 0.00 0.00 0.00 0.00 0.13 5.12
Self-Reminder 0.63 0.00 0.00 0.00 0.00 0.13 6.64
SmoothLLM 0.31 0.63 32.7 0.00 46.0 15.9 79.2
$\mathbb{IA}$ (Ours) 0.31 0.00 0.00 0.00 0.00 0.06 10.6
@@
Input Check 19.0 58.1 53.0 13.0 100 48.6 8.64
ICD 40.4 32.8 0.00 1.00 88.0 32.4 10.3
Self-Reminder 41.3 33.8 55.4 11.0 98.5 48.0 15.0
SmoothLLM 13.5 54.4 96.4 8.00 98.5 54.2 102.7
$\mathbb{IA}$ (Ours) 3.42 0.31 0.00 0.00 10.5 2.85 17.3
@@
Input Check 7.19 7.50 98.8 3.00 97.5 42.8 10.7
ICD 53.9 32.8 86.9 0.00 91.5 53.0 13.1
Self-Reminder 52.5 36.9 75.7 1.00 83.0 49.8 16.4
SmoothLLM 17.3 37.0 94.0 5.00 98.0 50.3 136.1
$\mathbb{IA}$ (Ours) 0.94 1.12 0.00 0.00 3.50 1.11 22.1
@@
Input Check 14.1 9.38 41.8 6.00 14.3 132.2
ICD 49.4 29.9 100.0 3.00 45.6 218.7
Self-Reminder 46.9 39.4 100 19.0 51.3 210.0
SmoothLLM 60.6 64.4 22.0 22.0 42.3 534.8
$\mathbb{IA}$ (Ours) 5.38 19.2 4.78 4.00 8.34 223.0
@@
Input Check 30.3 3.20 5.80 1.00 8.06 154.2
ICD 45.6 14.4 47.8 9.00 29.2 162.8
Self-Reminder 9.58 7.81 3.19 1.00 5.40 177.4
SmoothLLM 26.9 11.9 51.0 0.00 22.4 486.6
$\mathbb{IA}$ (Ours) 3.78 1.56 7.57 2.00 3.73 198.0
[…]
Input Check 2.50 0.00 0.00 4.00 1.63 2.47
ICD 0.94 0.31 0.00 0.00 0.31 5.12
Self-Reminder 2.81 0.31 0.00 0.00 0.78 7.21
SmoothLLM 0.64 0.00 0.00 0.00 0.16 15.2
$\mathbb{IA}$ (Ours) 0.64 0.00 0.00 0.00 0.16 8.27
Table 1: Comparison of our $\mathbb{IA}$ and four baselines under five jailbreak methods in terms of ASR (%) and time cost (s/sample). The best and second best average results are highlighted in bold and underline . Among them, DAN, SAP200, and DeepInception are complex and stealthy in-the-wild jailbreaks, while GCG and AutoDAN are optimization-based automatic jailbreaks. “—” means lacking official AutoDAN implementation for distributed larger models (MPT-30B and DeekSeek-67B) or white-box LLM weights required (GPT-3.5).